Please provide formula and process for solving 2.1 and 2.2 . NO EXCEL using please! Thank u!
Qustion2 Suppose the following information represents possible future returns for ATT stock next year: State of Economy Probability of state of Rate of return if state occurs 20 20 10% 20 1.00 (2.1) What is the expected rate of return for ATT common stock? (10 points) 22) What s the variance of the retum for ATT common stock? C0 pointy) ニ0. 2.10.1) + 04°-w、+1.25+ oos*) 0.2x(0.12-0
Answer #1

2.1. Expected Rate of Return of ATT Stock = ? Probability of Return* Expected Return

=0.2* 12% + 0.2* 6% - 0.2* 10% + 0.2* 22% + 0.2*5%

=7%

Q2. Variance= ? ((Mean Expected Return-Expected Return)^2)/Population Size

Mean Expected Return = (12+6-10+22+5)/5 =7%

Popultion Size=5

Therefore, Variance of ATT Common Stock Return = ((7-12)^2 + (7-6)^2 + (7-(-10))^2 +(7-22)^2+ (7-5)^2)/5

=108.8